# Slant sums of quiver gauge theories

### v1

@@ 1.6 Duals to zero-dimensional quiver varieties
In the same setting as the previous section, the 3d mirror dual to $\mathcal{M}$ is $\mathcal{M}^{!}=\overline{\mathcal{W}}^{\lambda}_{\mu}$, the corresponding affine Grassmannian slice. Combining (5) with Conjecture 1 from [16], and recalling that mirror symmetry identifies $\hbar^{!}=q/\hbar$, we obtain the following:
@@
The character of the tangent space of $\mathcal{M}^{!}$ at its unique torus fixed point $p^{!}$ is
$\sum_{\alpha\in\Phi^{-,\mathrm{re}}_{\mu}}\sum_{i=1}^{\langle\alpha,\mu\rangle}(\hbar^{!})^{1-i}e^{\alpha}+(\hbar^{!})^{i}e^{-\alpha},$
We show in Lemma 6.9 that $\sum_{\alpha\in\Phi^{-,\mathrm{re}}_{\mu}}\langle\alpha,\mu\rangle=\text{ht}(\lambda-\mu)=\sum_{i\in Q_{0}}\mathsf{v}_{i}$. The corresponding Coulomb branch $\mathcal{M}^{!}$ is birational to $T^{*}T_{\mathsf{v}}/W$, see [31], where $T_{\mathsf{v}}$ is a maximal torus of the gauge group $\prod_{i\in Q_{0}}GL(\mathsf{v}_{i})$ and $W$ is the Weyl group. Thus $\dim\mathcal{M}^{!}=2|\mathsf{v}|$, in agreement with Conjecture 1.5. This numerical coincidence provides some numerical explanation for the appearance of only real roots. It would be desirable to have a more conceptual explanation.
In Section 6, we will prove the following.
@@
Conjecture 1.5 holds in ADE type.
1.7 Partial resolutions of Coulomb branches
We expect that Conjecture 1.5 also provides a formula for the tangent space to an arbitrary partial resolution of a Coulomb branch at a nonsingular fixed point. In Section 5.4, we recall some facts about partial resolutions of Coulomb branches. We conjecture that isolated fixed points on the Higgs branch correspond to nonsingular torus fixed points on a partial resolution of the Coulomb branch. Furthermore, this correspondence works in such a way that the character of the tangent space at an arbitrary nonsingular fixed point on a partial resolution can be extracted from Conjecture 1.5, see Corollary 5.4.
@@ 1.8 Modules over quantized Coulomb branches
Assume that Conjecture 1.5 holds. We observe in Section 5.4 that the category $\mathcal{O}$ for a quantization of $\mathcal{M}^{!}$ contains a unique irreducible object and conjecture that its character can be read off from the explicit formula for the tangent space to the Coulomb branch at a fixed point (see Conjecture 5.7). More generally, we believe that Conjecture 1.5 should provide a formula for normalized characters of a natural family of modules over Coulomb branches labeled by nonsingular fixed points of their partial resolutions, see Remark 5.9 for a speculation about this.
@@ 1.9 Slant sums of Coulomb branches
In Section 7, we initiate the study of the slant sum construction from the Coulomb branch perspective. We prove that for $\mathsf{w}^{(2)}$ such that $\mathsf{w}^{(2)}_{\star_{2}}=1$ and all other $\mathsf{w}_{i}^{(2)}=0$, the slant sum on the Coulomb branch side corresponds to taking products (see Proposition 7.2). Note that the identification of a Coulomb branch $\mathcal{M}_{C}$ for a quiver $Q$ with the product $\mathcal{M}_{C}^{(1)}\times\mathcal{M}_{C}^{(2)}$ of Coulomb branches for $Q^{(1)},Q^{(2)}$ is compatible with natural structures that we have (integrable system, torus action) but only after a certain twist.
We also mention a partial description of the Coulomb branch for the slant sum for arbitrary $\mathsf{w}^{(2)}$ such that $\mathsf{w}^{(2)}_{\star_{2}}=1$ (see Proposition 7.7 that is an immediate corollary of [8, Proposition 3.18]). For $\mathsf{w}^{(2)}_{\star_{2}}>1$, we do not know how exactly $\mathcal{M}_{C}$, $\mathcal{M}_{C}^{(1)}$, $\mathcal{M}_{C}^{(2)}$ are related.
On the other hand, recall that the $\hbar=q$ specialized vertex functions restricted to fixed points are expected to equal graded traces of Verma modules over quantized Coulomb branches. So, our Higgs side computation suggests not a relation between (quantized) Coulomb branches for $Q,Q^{(1)}$, and $Q^{(2)}$ but between modules over some twisted versions of them.
Namely, one can consider the category of so-called Gelfand-Tsetlin modules over the quantized Coulomb branch $\mathcal{A}_{Q}$, see [42]. Let ${\mathsf{A}}_{\mathsf{v}}\subset G_{\mathsf{v}}$ be a maximal torus. It follows from [42] (see also [33, Appendix B] and [41] for the geometric treatment) that “blocks” of categories of GT-modules over $\mathcal{A}_{Q}$ are equivalent to categories of (topologically nilpotent) modules over the convolution algebra $\widehat{H}_{*}^{Z_{G}(\sigma_{\mathsf{v}})\times\mathsf{A}}(\mathcal{T}_{Q}^{(\sigma(t),t)}\times_{{\bf{N}}_{\mathcal{K}}^{(\sigma(t),t)}}\mathcal{T}_{Q}^{(\sigma(t),t)})$. Such a block depends on a choice of a cocharacter $\sigma\colon\mathbb{C}^{\times}\rightarrow\mathsf{A}_{\mathsf{v}}\times\mathsf{A}$. Here $\mathcal{T}_{Q}=G_{\mathcal{K}}\times^{G_{\mathcal{O}}}{\bf{N}}_{\mathcal{O}}$. Fixed points $\mathcal{T}_{Q}^{(\sigma(t),t)}$ can be explicitly computed and turn out to be equal to the disjoint unions of products of the corresponding fixed points on $\mathcal{T}_{Q^{(1)}}$ as well as fixed points on some twisted version of $\mathcal{T}_{Q^{(2)}}$, this is completely parallel to Theorem 3.1, where we provide a similar description of torus fixed based quasimaps to the resolved Higgs branch for $Q$. As a corollary of the above, one should be able to recover a branching formula for a graded trace of a GT-module over $\mathcal{A}_{Q}$ reflecting our Theorem 1.2 on the Coulomb side. We will return to this in the second version of the paper.
1.10 Branching for non-stationary Ruijsenaars function
In addition to providing a tool with which to study the vertex functions of zero-dimensional quiver varieties, Theorem 1.2 also subsumes and was motivated by an additional formula present in the literature, which we discuss in detail in Section 4.2. Let $X_{n}$ be the cotangent bundle to the variety of complete flags in $\mathbb{C}^{n}$. Viewed as a quiver variety, it decomposes as the slant sum of a cotangent bundle of a Grassmannian and a smaller flag variety: $X_{n}=T^{*}Gr(n-1,n)\#X_{n-1}$.
The vertex function of $X_{n}$ is the so-called “nonstationary Ruijsenaars function” studied, for example, in [9, 26, 35, 40]. It has the remarkable property of being solutions to the Macdonald difference equations in both the equivariant parameters for $\mathsf{A}$ and the Kähler variables.
Then formula (3.11) of [35] (see also (5.21)) is exactly Theorem 1.2 in this case. Furthermore, Theorem 1.2 also gives a similar branching formula for the cotangent bundles of partial flag varieties.
@@ 1.11 Outline
In Section 4, we consider some examples which show how Theorem 1.2 relates to Conjecture 1.3. We also explicitly spell out the connection between our results the branching of the non-stationary Ruijsenaars function of [35].
In Section 5, we recall some facts about Higgs and Coulomb branches and their mirror symmetry. We make some conjectures needed for Conjecture 1.5 to make sense. We also make some conjectures which would reduce the character of tangent spaces at fixed points for an arbitrary (partial resolution of a) Coulomb branch to the case of Conjecture 1.5.
In Section 6, we prove all of the conjectures in ADE types.
In Section 7, we investigate slant sums from the perspective of Coulomb branches.
@@ 1.12 Acknowledgements
We are grateful to Jiwu Jang, Shrawan Kumar, and Zijun Zhou for helpful conversations regarding this project. H. Dinkins was supported by NSF grant DMS-2303286 at MIT and the NSF RTG grant Algebraic Geometry and Representation Theory at Northeastern University DMS–1645877. R. Lance was supported by NSF grant DMS-2200867. Vasily Krylov was supported by the Simons Foundation Award 888988 as part of the Simons Collaboration on Global Categorical Symmetries.
@@ 2 Slant sums of Higgs branches
We briefly recall definitions, theorems, and fix conventions regarding quiver varieties. A more detailed introduction can be found in [32].
@@ 2.4 Split fixed points
Choosing a weight basis with respect to the ordering given by a $(p,i)$-chamber $\mathfrak{C}$, we obtain a canonical isomorphism $V_{i}\cong\mathbb{C}^{\mathsf{v}_{i}}$. Different choices of basis give isomorphisms that differ by a diagonal matrix in $GL(\mathbb{C}^{\mathsf{v}_{i}})$.
@@ 3.1 Review of quasimaps
As discussed in §2, Nakajima quiver varieties are defined as GIT quotients. so we may utilize the theory of quasimaps to a GIT quotient developed in [10]. In this paper, the domain of quasimaps will always be a parameterized $\mathbb{P}^{1}$.
@@
Let $\mathsf{QM}$ be the moduli space of stable quasimaps from $\mathbb{P}^{1}$ to $\mathcal{M}$. By [10], it is a Deligne-Mumford stack of finite type with a perfect deformation-obstruction theory. Thus it is equipped with a canonical virtual structure sheaf, which we will denote by $\mathcal{O}_{\mathrm{vir}}$. The canonical polarization $T^{1/2}$ of $\mathcal{M}$ pulls back under the universal evaluation morphism $\mathsf{QM}\times\mathbb{P}^{1}\to[\mu^{-1}(0)/G_{\mathsf{v}}]$, to a class $\mathscr{T}^{1/2}$, satisfying
@@
$\mathsf{QM}^{d}_{p}$$\mathsf{QM}^{d}_{\mathrm{ns}\,\infty}$$\{p\}$$\mathcal{M}$}
@@ 3.2 Twisted quasimaps
$\mathscr{W}^{\sigma}_{i}:=(\mathbb{C}^{2}\setminus\{0\})\times W)/\mathbb{C}^{\times}$
where $\mathbb{C}^{\times}$ acts by scaling on the first factor and on the second by $\sigma$. Equivalently, a quasimap from $\mathbb{P}^{1}$ to $\mathcal{M}$ is a section of the trivial $\mu^{-1}(0)/G_{\mathsf{v}}]$ bundle on $\mathbb{P}^{1}$; a $\sigma$-twisted quasimap is a section of a nontrivial $\mu^{-1}(0)/G_{\mathsf{v}}]$ bundle on $\mathbb{P}^{1}$ whose topology is determined by $\sigma$.
@@ 3.3 Vertex functions
The restriction of the evaluation map $\operatorname{ev}_{\infty}:\mathsf{QM}^{d}_{\mathrm{ns}\,\infty}\to\mathcal{M}$ to the $\mathbb{C}^{\times}_{q}$-fixed locus is proper, see [37, §7.3]. So we obtain a pushforward in localized equivariant $K$-theory,
@@
The symmetrized virtual structure sheaf is defined by
$\hat{\mathcal{O}}_{\operatorname{vir}}=\mathcal{O}_{\mathrm{vir}}\otimes\left(\mathscr{K}_{\mathrm{vir}}\frac{\det\mathscr{T}^{1/2}|_{\infty}}{\det\mathscr{T}^{1/2}|_{0}}\right)^{1/2}$ (10)
where $\mathscr{K}_{\mathrm{vir}}=\operatorname{ev}_{\infty}^{*}(\det T\mathcal{M})\otimes\left(\det\mathscr{T}_{\mathrm{vir}}\right)^{-1}$ is the (normalized) virtual canonical bundle. The existence of the square root is discussed in [37, §6.1.8].
@@
$V^{(\tau)}(z)=\sum_{d\in\operatorname{Eff}(\mathcal{M})}(\operatorname{ev}_{\infty})_{*}\left(\mathsf{QM}^{d}_{\mathrm{ns}\,\infty},\operatorname{ev}_{0}^{*}(\tau)\otimes\hat{\mathcal{O}}_{\operatorname{vir}}\right)z^{d}$
@@
The case of $\tau=1$ is sometimes referred to as the bare vertex function and denoted $V(z):=V^{(1)}(z)$. For $p\in\mathcal{M}$, we denote the restriction by $V^{(\tau)}_{p}(z):=V^{(\tau)}(z)$.
@@ 3.4 Branching rule for vertex functions
There are now several quiver varieties present and we will consider based quasimaps to each of them. The notation $\mathsf{QM}_{p}$ and $\mathsf{QM}_{p^{(1)}}$ now refer to quasimaps to the different varieties $\mathcal{M}$ and $\mathcal{M}^{(1)}$, respectively, rather than quasimaps to a fixed variety based at different points.
@@
where $F$ runs over all connected components of $\left(\mathsf{QM}^{d^{(1)}}_{p^{(1)}}\right)^{\mathsf{T}\times\mathbb{C}^{\times}_{q}}$. Furthermore, this isomorphism identifies the virtual tangent spaces as $\mathsf{T}$, where the virtual tangent spaces on the right hand side are pulled back via $\iota$.
@@
Since it identifies the virtual tangent spaces, the isomorphism above respects the virtual structure sheaves. Since we assume the canonical polarization is used in (10), it follows that the isomorphism respects the symmetrized virtual structure sheaves as well. Recall the map $\iota$ from (1) and let $\tau=\iota^{*}(\tau_{1}\otimes\tau_{2})$. We write the Kähler parameters similarly as $z=(z_{1},z_{2})$.
@@
where $T^{\sigma}_{p}\mathcal{M}$ is the tangent space of $\mathcal{M}$ at $p$, viewed as a $\mathsf{T}\times\mathbb{C}^{\times}_{q}$ representation where the second factor acts by $\sigma$. As in [37, §8.2], this leads to a relationship between the twisted and untwisted vertex functions.

### v2

@@ 1.6 Duals to zero-dimensional quiver varieties
Recall that $\mathcal{M}$ is a single point if and only if $\mu\in W\lambda$. Combining (4) with Conjecture 1 from [undefr], and recalling that mirror symmetry identifies $\hbar^{!}=q/\hbar$, we obtain the following:
@@
If $\mu\in W\lambda$, then the unique torus fixed point $p^{!}\in\mathcal{M}^{!}$ is nonsingular, and the character of the tangent space of $\mathcal{M}^{!}$ at $p^{!}$ is
$\sum_{\alpha\in\Phi^{-,\mathrm{re}}_{\mu}}\sum_{i=1}^{\langle\alpha,\mu\rangle}(\hbar^{!})^{1-i}e^{\alpha}+(\hbar^{!})^{i}e^{-\alpha}.$ (5)
We provide some numerical evidence for Conjecture 1.7 in Lemma 7.10.
By [undefap], the Coulomb branch is well-defined for quivers encoding symmetrizable Cartan matrices. We thus see that Conjecture 1.7 makes sense more generally for any quiver $Q$ such that $\mathfrak{g}_{Q}$ is symmetrizable. Here it becomes crucial that we treat $\mu$ as a coweight (instead of a weight as is often the case) and the $\alpha$ as roots in (5), equipped with the natural pairing between them.
Using the identification of Coulomb branches for finite Dynkin quivers with affine Grassmannian slices, we will prove in Section 7 the following.
@@
Conjecture 1.7 holds for finite type Dynkin quivers.
Our argument for the proof of Proposition 1.8 is obtained by combining the argument similar to the one in [undefaf, Lemma 4.4] with [undefah, Theorem 3.1(1)].
@@ 1.7 Modules over quantized Coulomb branches
Assume that Conjecture 1.7 holds. In Section 6, we observe that the category $\mathcal{O}$ for a quantization of $\mathcal{M}^{!}$ contains a unique irreducible object, and we show that its character can be read off from the explicit formula for the tangent space to the Coulomb branch at a fixed point (see Proposition 6.11). Our main technical tool is the $D$-module of graded traces introduced in [undefad] and further studied in [undefq].
Restricting to finite-type Dynkin quivers, and using that Conjecture 1.7 holds in this case (see Proposition 1.8 above), we obtain explicit formulas for the characters of so-called “extremal” (terminology is due to [undefab]) irreducible modules $L$ over shifted Yangians $Y_{\mu}$. Namely, we get (see Proposition 6.11):
$\widetilde{\operatorname{ch}}(L)=\prod_{\alpha\in\Phi_{\mu}^{-}}\frac{1}{(1-e^{\alpha})^{\langle\alpha,\mu\rangle}},$ (6)
where $\widetilde{\operatorname{ch}}(L)$ is the normalized character of $L$.
Using [undefay, Corollary 1.2.1], these formulas may be translated to the case of shifted quantum groups. A closely related result was obtained by Negut by completely different methods (see [undefaq, Theorem 1.3]). The authors of the forthcoming paper [undefab] also obtain the same formula for the extremal $Y_{\mu}$-modules; their argument combines [undefah, Theorem 3.1(1)] with the theory developed in [undefak].
Finally, the character formula (6) we obtain admits a natural one-parameter deformation coming from the parameter $\hbar^{!}$ in (5) above. This refined character agrees with that in [undefaq, Theorem 1.3]. For us, the parameter $\hbar^{!}$ corresponds to the $\mathbb{C}^{\times}$-action on $\mathcal{M}^{!}$ induced by the homological grading on $\mathbb{C}[\mathcal{M}^{!}]$.
1.8 Connection with quantum Hikita conjecture
Corollary 1.4 is related to another manifestation of 3d mirror symmetry: the quantum Hikita conjecture proposed in [undefad]. This conjecture predicts an isomorphism between the $q=\hbar$ specialized quantum $D$-module for the Higgs branch and the $D$-module of graded traces for the Coulomb branch.
A proof of this conjecture was given for ADE quiver varieties with minuscule framings in [undefq]. The technical heart of the proof is to identify the $q=\hbar$ specialized descendant vertex functions with the graded traces of elements of the quantized Coulomb branch–in this case a truncated shifted Yangian–over a Verma module. Thus Corollary 1.4 provides many cases of the Higgs side of the computation needed to prove the quantum Hikita for more general quiver varieties. In many examples, see Section 4.1, vertex functions can be computed in terms of reverse plane partitions over some poset. Theorem 1.2 shows that this is true for $V^{(\tau)}_{p}$ if it is true for $V^{(\tau_{1})}_{p^{(1)}}$ and $V^{(\tau_{2})}_{p^{(2)}}$.
Furthermore, passing Corollary 1.4, or more generally the $q=\hbar$ specialization of Theorem 1.2, through 3d mirror symmetry gives a conjectural “branching rule” for graded traces. To the best of the authors’ knowledge, this formula, and its expression in terms of reverse plane partitions in special cases, is a new expectation.
@@ 1.9 Slant sums of Coulomb branches
In Section 8, we initiate the study of the slant sum construction from the Coulomb branch perspective. We prove in Proposition 8.3 that for $\mathsf{w}^{(2)}$ such that $\mathsf{w}^{(2)}_{\star_{2}}=1$ and all other $\mathsf{w}_{i}^{(2)}=0$, the slant sum on the Coulomb branch side corresponds to taking products. The Higgs branch counterpart of this statement was observed in [undefb, Lemma 8.4] (after applying the Crawley-Boevey trick). We note that the identification of a Coulomb branch $\mathcal{M}_{C}$ for the quiver $Q=Q^{(1)}{}_{\star_{1}}\#_{\star_{2}}Q^{(2)}$ with the product $\mathcal{M}_{C}^{(1)}\times\mathcal{M}_{C}^{(2)}$ of Coulomb branches for $Q^{(1)},Q^{(2)}$ is compatible with natural structures (integrable system, torus action) after a certain twist.
@@ 1.10 Outline
In Section 4, we consider some examples which show how Theorem 1.2 relates to Conjecture 1.5. We also explicitly spell out the connection between our results the branching of the non-stationary Ruijsenaars function of [undefas].
In Section 5, we recall some facts about Higgs and Coulomb branches and their mirror symmetry. We make some conjectures needed for Conjecture 1.7 to make sense. We also make some conjectures which reduce the character of tangent spaces at nonsingular fixed points for an arbitrary partial resolution of a Coulomb branch to the case of Conjecture 1.7.
In Section 6 we prove a general result describing the normalized character of irreducible modules over certain Coulomb branches in presence of Conjecture 1.7 above. In particular, we obtain a refined character formulas for so-called “extremal” irreducible modules over shifted Yangians.
In Section 7, we prove all of the conjectures above for finite type Dynkin quivers.
Finally, in Section 8, we investigate slant sums from the perspective of Coulomb branches.
@@ 1.11 Acknowledgements
We are grateful to Jiwu Jang, Shrawan Kumar and Zijun Zhou for helpful conversations regarding this project. We are grateful to Ivan Losev and Dmytro Matvieievskyi for helpful discussions on the material of Section 6. We are grateful to Joel Kamnitzer, Alexis Leroux-Lapierre, and Hiraku Nakajima for valuable comments on an earlier version of this text. H. Dinkins was supported by NSF grant DMS-2303286 at MIT and the NSF RTG grant Algebraic Geometry and Representation Theory at Northeastern University DMS–1645877. R. Lance was supported by NSF grant DMS-2200867. Vasily Krylov was supported by the Simons Foundation Award 888988 as part of the Simons Collaboration on Global Categorical Symmetries.
@@ 2 Slant sums of Higgs branches
We briefly recall definitions, theorems, and fix conventions regarding quiver varieties. A more detailed introduction can be found in [undefao].
@@ 2.4 Split fixed points
Choosing a weight basis with respect to the ordering given by a $(p,i)$-chamber $\mathfrak{C}$, we obtain an isomorphism $V_{i}\cong\mathbb{C}^{\mathsf{v}_{i}}$. Different choices of basis give isomorphisms that differ by a diagonal matrix in $GL(\mathbb{C}^{\mathsf{v}_{i}})$.
@@ 3.1 Review of quasimaps
As discussed in §2, Nakajima quiver varieties are defined as GIT quotients. so we may utilize the theory of quasimaps to a GIT quotient developed in [undefl]. In this paper, the domain of quasimaps will always be a parameterized $\mathbb{P}^{1}$.
@@
Let $\mathsf{QM}$ be the moduli space of stable quasimaps from $\mathbb{P}^{1}$ to $\mathcal{M}$. By [undefl], it is a Deligne-Mumford stack of finite type with a perfect deformation-obstruction theory. Thus it is equipped with a canonical virtual structure sheaf, which we will denote by $\mathcal{O}_{\mathrm{vir}}$. The canonical polarization $T^{1/2}$ of $\mathcal{M}$ pulls back under the universal evaluation morphism $\mathsf{QM}\times\mathbb{P}^{1}\to[\mu^{-1}(0)/G_{\mathsf{v}}]$, to a class $\mathscr{T}^{1/2}$, satisfying
@@
$\mathsf{QM}^{d}_{p}$$\mathsf{QM}^{d}_{\mathrm{ns}\,\infty}$$\{p\}$$\mathcal{M}$$\scriptstyle{\operatorname{ev}_{\infty}}$
@@ 3.2 Twisted quasimaps
$\mathscr{W}^{\sigma}_{i}:=(\mathbb{C}^{2}\setminus\{0\})\times W_{i})/\mathbb{C}^{\times}$
where $\mathbb{C}^{\times}$ acts by scaling on the first factor and on the second by $\sigma$.
@@ 3.3 Vertex functions
The restriction of the evaluation map $\operatorname{ev}_{\infty}:\mathsf{QM}^{d}_{\mathrm{ns}\,\infty}\to\mathcal{M}$ to the $\mathbb{C}^{\times}_{q}$-fixed locus is proper, see [undefau, §7.3]. So we obtain a pushforward in localized equivariant $K$-theory,
@@
The symmetrized virtual structure sheaf on $\mathsf{QM}_{\mathrm{ns}\,\infty}^{d}$ is defined by
$\hat{\mathcal{O}}_{\operatorname{vir}}^{d}=\mathcal{O}_{\mathrm{vir}}^{d}\otimes\left(\mathscr{K}_{\mathrm{vir}}\frac{\det\mathscr{T}^{1/2}|_{\infty}}{\det\mathscr{T}^{1/2}|_{0}}\right)^{1/2}$ (11)
where $\mathscr{K}_{\mathrm{vir}}=\operatorname{ev}_{\infty}^{*}(\det T\mathcal{M})\otimes\left(\det\mathscr{T}_{\mathrm{vir}}\right)^{-1}$ is the (normalized) virtual canonical bundle. The existence of the square root is discussed in [undefau, §6.1.8].
@@
$V^{(\tau)}(z)=\sum_{d\in\operatorname{Eff}(\mathcal{M})}(\operatorname{ev}_{\infty})_{*}\left(\mathsf{QM}^{d}_{\mathrm{ns}\,\infty},\operatorname{ev}_{0}^{*}(\tau)\otimes\hat{\mathcal{O}}_{\operatorname{vir}}^{d}\right)z^{d}$
@@
The case of $\tau=1$ is sometimes referred to as the bare vertex function and denoted $V(z):=V^{(1)}(z)$. For $p\in\mathcal{M}^{\mathsf{T}}$, we denote the restriction by $V^{(\tau)}_{p}(z):=V^{(\tau)}(z)$.
@@ 3.4 Branching rule for vertex functions
There are now several quiver varieties present and we will consider based quasimaps to each of them. The notation $\mathsf{QM}_{p}$ and $\mathsf{QM}_{p^{(1)}}$ now refer to based quasimaps to the different varieties $\mathcal{M}$ and $\mathcal{M}^{(1)}$, respectively, rather than quasimaps to a fixed variety based at different points.
@@
where $F$ runs over all connected components of $\left(\mathsf{QM}^{d^{(1)}}_{p^{(1)}}\right)^{\mathsf{T}\times\mathbb{C}^{\times}_{q}}$. Furthermore, this isomorphism identifies the virtual tangent spaces as $\mathsf{T}$ representations, where the virtual tangent spaces on the right hand side are pulled back via $\iota$.
@@
Since it identifies the virtual tangent spaces, the isomorphism above respects the virtual structure sheaves. Since we assume the canonical polarization is used in (11) for all the symmetrized virtual sheaves, it follows that the isomorphism respects the symmetrized virtual structure sheaves as well. Recall the map $\iota$ from (1) and let $\tau=\iota^{*}(\tau_{1}\otimes\tau_{2})$. We write the Kähler parameters similarly as $z=(z_{1},z_{2})$.
@@
where $T^{\sigma}_{p}\mathcal{M}$ is the tangent space of $\mathcal{M}$ at $p$, viewed as a $\mathsf{T}\times\mathbb{C}^{\times}_{q}$ representation where the second factor acts by $\sigma$. As in [undefau, §8.2], this leads to a relationship between the twisted and untwisted vertex functions.